OpenStudy (anonymous):

-3a is always, sometimes, or never a negative value. Explain.

TheSmartOne (thesmartone):

if a is negative then what is -3a? if a is positive then what is -3a?

OpenStudy (igreen):

a could be anything..it could be positive or negative. A positive times a negative = a negative A negative times a negative = a positive

OpenStudy (igreen):

@heatherg75 What do you think the answer is?

OpenStudy (igreen):

@logancoop Don't give out answers

OpenStudy (anonymous):

No direct answers. @logancoop

OpenStudy (anonymous):

Oops

TheSmartOne (thesmartone):

Delete yuor reply.

OpenStudy (anonymous):

Welcome to OpenStudy, good job on your first question! :)

OpenStudy (hauntedwoodsgal):

If you mean: −|A|+2∗B<0 Then it is sometimes true, some examples: A = 1, B = 2 −(1)+2∗(2)=3 which of course is not negative, so in this case is false or A = 5, B = 1 −(5)+2∗(1)=−3 which is negative and in this case is true.

TheSmartOne (thesmartone):

@heatherg75 have you gotten an answer?